Exercise 7-6 Variable and Absorption Costing Unit Product Costs and Income Statements [LO7-1, LO7-2]

Lynch Company manufactures and sells a single product. The following costs were incurred during the company’s first year of operations:

Variable costs per unit:
Manufacturing:
Direct materials $ 12
Direct labor $ 4
Variable manufacturing overhead $ 2
Variable selling and administrative $ 2
Fixed costs per year:
Fixed manufacturing overhead $ 370,000
Fixed selling and administrative $ 200,000

During the year, the company produced 37,000 units and sold 18,000 units. The selling price of the company’s product is $55 per unit.

Required:

1. Assume that the company uses absorption costing:

a. Compute the unit product cost.

b. Prepare an income statement for the year.

2. Assume that the company uses variable costing:

a. Compute the unit product cost.

b. Prepare an income statement for the year.
Answer #1
1a
Direct materials 12
Direct labor 4
Variable manufacturing overhead 2
Fixed manufacturing overhead 10 =370000/37000
Unit product cost 28
b
Absorption Costing Income Statement
Sales 990000 =18000*55
Cost of goods sold 504000 =18000*28
Gross margin 486000
Selling and administrative expenses 236000 =200000+(18000*2)
Net operating income (loss) 250000
2a
Direct materials 12
Direct labor 4
Variable manufacturing overhead 2
Unit product cost 18
b
Variable Costing Income Statement
Sales 990000 =18000*55
Variable expenses:
Variable Cost of goods sold 324000 =18000*18
Variable Selling and administrative expenses 36000 =18000*2
360000
Contribution margin 630000
Fixed expenses:
Fixed manufacturing overhead 370000
Fixed Selling and administrative expenses 200000
570000
Net operating income (loss) 60000